### v1

# Bose-Hubbard model with power-law hopping in one dimension

Abstract
We investigate the zero-temperature phase diagram of the one-dimensional Bose-Hubbard model with power-law hopping decaying with distance as $1/r^{\alpha}$ using exact large scale Quantum Monte-Carlo simulations. For all $1<\alpha\leq 3$ the quantum phase transition from a superfluid and a Mott insulator at unit filling is found to be continuous and scale invariant, in a way incompatible with the Berezinskii-Kosterlitz-Thouless (BKT) scenario, which is recovered for $\alpha>3$. We characterise the new universality class by providing the critical exponents by means of data collapse analysis near the critical point for each $\alpha$ and from careful analysis of the spectrum. Large-scale simulations of the grand canonical phase diagram and of the decay of correlation functions demonstrate an overall behavior akin to higher dimensional systems with long-range order in the ground state for $\alpha\leq 2$ and intermediate between one and higher dimensions for $2<\alpha\leq 3$. Our exact numerical results provide a benchmark to compare theories of long-range quantum models and are relevant for experiments with cold neutral atom, molecules and ion chains.
The Bose-Hubbard (BH) model describes the dynamics of interacting bosons confined in a lattice potential with nearest-neighbor hopping energy $t$ and local interactions $U$. It features a localization quantum phase transition from a gapless superfluid (SF) to a gapped Mott insulator (MI) as a function of the ratio $t/U$ [1, 2, 3, 4]. The BH model has been successfully used to describe quantum phase transitions in systems as diverse as cold atoms trapped in optical lattices, superfluid ${}^{4}{\rm He}$, and superconductors. In one dimension, the BH quantum phase transition at constant integer density $\rho$ belongs to the Berezinskii-Kosterlitz-Thouless (BKT) [5, 6, 7, 8, 9, 10] universality class. The latter has been experimentally demonstrated in cold atom experiments in Refs. [11, 12] and is widely believed to underpin all localization transitions in one dimension.
In the last few years, advances in engineering Hamiltonians with Rydberg atoms, cold dipolar atoms and molecules, trapped ions coupled to motional degrees of freedom and neutral atoms coupled to photonic modes [13, 14, 15, 16, 17] have sparked significant interest in the many-body physics of quantum models with long-range couplings. Theory and experiments have provided evidence for novel static and dynamic phenomena in these systems [18], such as, e.g., the non-local propagation of correlations [19, 20, 21, 22, 23], breaking of conformal symmetry [24, 25], new topological phases of matter and phase transitions [24, 26, 27]. The most interesting regime is that of so-called “weak long-range interactions” with $d<\alpha<\alpha_{*}$, which is intermediate between the limit of infinite-range “strong interactions” for $\alpha<d$ and short-range-like interactions for $\alpha>\alpha_{*}$, with $d$ the dimension and $\alpha_{*}$ a threshold value that depends on the system and transition under study [18]. While integrable models provide a guidance in some situations [24] and despite an intense theoretical effort, it remains an open challenge to precisely characterize quantum phases and phase transitions in non-integrable quantum models with long-range couplings. It is thus of fundamental importance to obtain exact results for these systems, to which theories can be compared.
In this work, we investigate the phase diagram of the $1d$ BH model with power-law hopping that decays with distance as $1/r^{\alpha}$ in the regime of weak long-range couplings $\alpha>1$. We use exact large-scale Quantum Monte Carlo simulations based on the Worm Algorithm [28] to determine the ground state phase diagram and to characterize the superfluid phases. Qualitatively, the phase diagram in the grand-canonical ensemble shows a shrinking of the MI lobes in the $t/U$-$\mu/U$ plane ($\mu$ is the chemical potential) with respect to short-range models for all $\alpha<3$ [29] and a rounding of the lobes akin to higher-dimensional models for $\alpha<2$. Interestingly, for constant $\rho=1$ a finite-size scaling analysis of the winding number fluctuations, measuring superfluid properties, shows that the superfluid-Mott insulator quantum phase transition is incompatible with the BKT universality class for any $1<\alpha\leq 3$. It corresponds instead to a new continuous scale-invariant phase transition in the whole parameter range of weak long-range couplings. We characterize this new universality class for bosons in one dimension by determining the critical exponents by data collapse at the critical point and the energy spectrum. For $\alpha>3$ the system resembles short-range models, which fixes $\alpha_{*}$ to $\alpha_{*}=3$. From a large-scale analysis of correlation functions, we confirm that long-range order exists for $\alpha<2$ in the superfluid phase, in agreement with literature, while we find no evidence of such order for $\alpha>2$. These exact results provide benchmarks for theories and experiments.
The $1d$ BH model with power-law hopping reads
$\mathcal{H}=-t\sum_{i<j}\frac{a^{\alpha}}{|r_{ij}|^{\alpha}}\left[{b}^{\dagger}_{i}{b}_{j}+\text{H.c.}\right]+\frac{U}{2}\sum_{i}{n}_{i}({n}_{i}-1)-\mu\sum_{i}{n}_{i}$ (1)
Here, ${b}^{\dagger}_{i}$, ${b}_{i}$, and ${n}_{i}={b}^{\dagger}_{i}{b}_{i}$ are the bosonic creation, annihilation and particle number operators on site $i$, respectively; $t$, $U$ and $\mu$ are the hopping energy, the on-site interaction energy and the chemical potential, respectively, with $a$ the lattice spacing. The energy and length scales are set by choosing $t=1$ and $a=1$. For nearest neighbor hopping ($\alpha\rightarrow\infty$) and density $\rho=1$, Eq. (1) displays a zero-temperature quantum phase transition of the BKT type from a superfluid to a Mott insulator at a critical value $(t/U)_{c}=0.300\pm 0.025$ [30, 31, 32, 33]. For finite $\alpha$ and $U/t\rightarrow\infty$ (hard-core bosons), Eq. (1) maps into a long-range XY model, for which spin-wave and semi-analytical renormalization group analyses predict a continuously varying dynamical exponent $z=(\alpha-1)/2$ for $\alpha<3$ [34] as well as a breaking of $U(1)$ symmetry, and ensuing long-range order, for a given $\alpha_{c}<3$, to be determined numerically. In Ref. [35], $\alpha_{c}$ was estimated to be $\alpha_{c}\simeq 2.8$ using a density matrix renormalization group approach for system sizes up to $L\simeq 100$. In this work, we investigate Eq. (1) for $1<\alpha\leq 3$ and all values $t/U$ via large scale quantum Monte Carlo simulations of up to $L=1024$ sites and inverse temperature $\beta=L^{z_{*}}$, small enough to probe ground state properties. We focus first on the superfluid-Mott insulator quantum phase transition at varying and constant densities and then discuss the correlation functions in the liquid phase.
Figure 1: Ground state phase diagram of the 1D Bose-Hubbard model with power-law hopping for $\alpha=1.6$, $1.9$, $2.5$, $2.7$, and $3.0$, showing the boundary between the Mott insulator (at unity filling) and superfluid phases as a function of chemical potential $\mu/U$ and hopping amplitude $t/U$. Markers represent simulation data, while the solid black line (for $\alpha=1.6$) and the dashed black line (for $\alpha=1.9$) are obtained from fitting the energy gap $\Delta/t$ near the critical point as $\Delta/t\sim|t/U-{t/U}_{\rm c}|^{-z_{*}\nu}$ (see also Fig. 2 and text). Inset: Gap Energy $\Delta/t$ fitted as a function of $x\equiv t/U-{t/U}_{\rm c}$.
Figure 2: Dispersion relation $E(k)$ vs. $k$ for $\alpha=1.6$, $1.9$, $2.5$, and $2.7$ in the superfluid phase near the transition along the commensurate density line. Symbols denote numerical data obtained from fitting the long-time exponential decay of $G(k,\tau)$, while lines show the spin-wave analysis prediction $E(k)\sim k^{z_{*}}$, with $z_{*}=(\alpha-1)/2$ [34]. Inset: Single-particle Green function $G(k,\tau)$ for $k=\pi/64$ (represented by empty markers) and $\alpha=1.6,1.9,2.5,$ and $2.7$ for $L=256$. The black dashed line represents the numerical fit to the exponential decay.
Figure 1 shows the ground state phase diagram as a function of $\mu/U$ and $t/U$ for different $\alpha\leq 3$. For each $\alpha$, the figure shows the existence of a lobe, corresponding to a MI phase at unit filling surrounded by a SF phase. For each ratio $t/U$, the boundaries in $\mu$ of the lobe are determined by computing the energy gap in the MI phase from the Green function $G(k=0,\tau)$, which is obtained via the spatial averaging of the Matsubara Green function, $G(i,\tau)=\langle b_{i}^{\dagger}(\tau)b_{0}(0)\rangle$, with $\tau\in[-\beta/2,\beta/2]$ the imaginary time and $k$ the quasi momentum. By employing the Lehmann expansion, the Green function behaves as $G(k,\tau)\propto e^{-\epsilon_{\pm}(k)|\tau|}$ for $\tau\rightarrow\pm\infty$, where ϵ±subscriptitalic-$\epsilon_{\pm}$ denote the particle and hole energies, respectively [36, 37] (see also example in Inset of Fig. 2 and below): For $\tau>0$, $G(k,\tau)$ describes a particle excitation, whereas for $\tau<0$, it represents a hole excitation in the MI phase. By fitting the asymptotic behavior of $G(k=0,\tau)$ to an exponential form, the particle and hole excitation energies are extracted from the slopes. These excitation energies are measured relative to the chemical potential in the grand canonical ensemble. The insulating gap, $\Delta$, is then determined as $\Delta=\mu_{+}-\mu_{-}$, where $\mu_{\pm}=\mu\pm\epsilon_{\pm}$ and $\mu$ is the chemical potential used in the simulation.
Figure 1 shows two distinct behaviours of the lobe structure for $\alpha<2$ and $\alpha>2$. Surprisingly, for $\alpha<2$ the Mott lobes are smooth and rounded, similar to those observed for corresponding MI phases in higher dimensions [38, 37]. For $\alpha>2$, instead, the MI phase exhibits a progressively more pointed and asymmetric structure with increasing $\alpha$ and the critical point at constant density (i.e. the tip of the former lobe) shifts to larger values of $t/U$. This behavior is reminiscent of the needle-type shape observed for the short-range hopping model with the BKT transition at $\rho=1$, indicating a behavior intermediate between higher dimensions and one dimension. We come back below to the precise nature of the critical point.
We further investigate the properties of the superfluid phase near the transition point at constant density by computing the excitation spectrum. Figure 2 shows the dispersion relation $E(k)$ vs $k$ for different $31<\alpha<$ 3, where quantum Monte-Carlo data for $E(k)$ (symbols) are obtained by fitting the large-$\tau$ decay of the Green’s function $G(k,\tau)$ to an exponential form (see Inset for examples). The resulting dispersion relations are sub-linear in $k$, in excellent agreement with the predicted scaling $E(k)\sim k^{z_{*}}$ from spin-wave theory for all $\alpha$, with $z_{*}=(\alpha-1)/2$ [34].
Figure 3: Characterization of the Mott insulator to superfluid phase transition: (a)-(b) Mean-square winding number $\langle\mathcal{W}^{2}\rangle$ vs $t/U$ for $\alpha=1.6$ and $2.7$ near phase transitions, showing crossing between the curves; the curve corresponding to the largest lattice size is subtracted from all data for clarity. Vertical error bars indicate the estimated uncertainty from the Monte Carlo simulations. Insets: Finite-size scaling of crossings points between curves for system sizes $L_{1}$ and $L_{2}=2L_{1}$ as a function of $L=L_{1}$. (c) Phase diagram, ${t/U}_{c}$ vs $\alpha$ of the Mott insulator and superfluid quantum phases for model (1).
Figure 4: Superfluid stiffness $Y_{s}$ near the critical interaction strength ${t/U}_{c}$ for $\alpha=1.6,1.9$ and $2.7$ and lattice size $L=128,256,512$. Inset: Data collapse for the scaled superfluid stiffness $L^{-\zeta/\nu}Y_{s}$ vs $L^{1/\nu}[(t/U)-({t/U}_{c}]$. The fitted values of the correlation length exponent $\nu$ are directly reported in the inset. For $\alpha=1.6$ and $1.9$, $\nu$ is consistent with the value derived by fitting energy gap $\Delta$. (see Figure 1)
Figure 5: Characterization of the superfluid phase: (a)-(d) Single-particle density matrix, $\mathcal{G}(\ell)$, plotted against $\ell$ for $\alpha=1.3$, $1.6$, $2.5$, and $2.7$. The dashed line represents the best fit to $A\cdot c(\ell)^{-\gamma}$, where $c(\ell)=\sin(\pi\ell/L)$ is the chord distance, and $A$ and $\gamma$ are fitting parameters. For $\alpha=1.3$ and $1.9$, $\mathcal{G}(\ell)$ saturates to a constant as $\ell\rightarrow\infty$. For $\alpha=2.5$ and $2.7$, $\mathcal{G}(\ell)$ exhibits algebraic decay. Insets: Difference $\Delta\mathcal{G}(\ell)$ between $\mathcal{G}(\ell)$ for system size $L$ and the numerical fit for $L=L_{\text{max}}$, plotted as a function of the chord distance $c(\ell)$. The difference indicates an upward trend of $\mathcal{G}(\ell)$ for increasing lattice size. (e) Finite-size scaling of the power-law exponent $\gamma$ to the thermodynamic limit.
We characterize the SF-MI quantum phase transition at constant density $\rho=1$ by computing the mean-squared winding number fluctuations $\langle\mathcal{W}^{2}\rangle$ – a scale-invariant quantity proportional to the superfluid stiffness $Y_{s}$ as $\langle\mathcal{W}^{2}\rangle=Y_{s}/(LT)$ – up to sizes $L=512$. $\langle\mathcal{W}^{2}\rangle$ is expected to have a finite value and a zero value in the SF and MI phases, respectively. Figures 3(a) and (b) present example results for $\langle\mathcal{W}^{2}\rangle$ as a function of $t/U$ for two power-law exponents $\alpha<2$ and $\alpha>2$ in Eq. (1), respectively, and for different system sizes $L$. In the figures, the values of $\langle\mathcal{W}^{2}\rangle_{L_{\text{max}}}$ for the largest sizes $L_{\text{max}}$ used in the computations have been subtracted for clarity. Panels (a) and (b) show a clear crossing of $\langle\mathcal{W}^{2}\rangle$ vs $t/U$ when plotted for different $L$ at values $(t/U)_{c}=0.043\pm 0.0005$ and $(t/U)_{c}=0.131\pm 0.0008$, respectively. These crossings correspond to a quantum phase transition at the respective values of the critical ratio $(t/U)_{c}$ (see Insets for further extrapolation to thermodynamic limit). Interestingly, the very presence of a crossing in the $\langle\mathcal{W}^{2}\rangle-(t/U)$ curves rules out the Berezinskii-Kosterlitz-Thouless universality class for these power law models, in contrast to familiar short-range models in one dimension [10] and long-range models with power-law density-density interactions [39, 40, 10]. Similar crossings are observed for all $1<\alpha\leq 3$, implying a continuous scale-invariant phase transition in this whole range of $\alpha$. For $\alpha>3$ the transition is instead of the BKT type, consistent with short-range models. This fixes $\alpha_{*}=3$ for the model (1).
Figure 3(c) summarizes the ground state phase diagram of Eq. (1) as a function of $(t/U)_{c}$ and $1<\alpha\leq 3$. In this diagram, $(t/U)_{c}$ for each value of $\alpha$ is identified based on the scale-invariant crossing point, as previously detailed. The discovery of this family of scale-invariant phase transitions is a key result of this work.
We further characterize the SF-MI quantum phase transitions at commensurate density by determining the correlation length exponent $\nu$ associated to the continuous transition using data collapse analysis near the critical points. For each $\alpha$, we rescale the superfluid stiffness using $L^{-\zeta/\nu}Y_{s}$ as a function of $L^{1/\nu}[(t/U)-(t/U)_{c}]$ , where $\nu$ and $\zeta$ are fitting parameters (see Insets). These parameters are determined through optimization using the Nelder-Mead algorithm [41], with a cost function based on the Kawashima-Ito-Houdayer-Hartmann quality metric [42, 43]. Example results for $\alpha=1.6,1.9,$ and $2.7$ are shown in Figure 4, demonstrating good data collapse near the critical points for all $\alpha\geq 1$ using power-law rescaling, which further proves the scale-invariant nature of the phase transition.
For consistency, we further estimate the correlation length exponent $\nu$ by using the expected expression for the gap energy $\Delta/t\sim|t/U-{t/U}_{\rm c}|^{-z_{*}\nu}$ near the critical point (see Fig. 1), with $z_{*}$ obtained numerically from the energy dispersion relation, as discussed above. As an example, for $\alpha=1.6$ and 1.9 we obtain $\nu=1.77\pm 0.03$ and $\nu=1.49\pm 0.04$, respectively, in good agreement with the estimate from data collapse. Our results for different $\alpha$ values show that $\nu$ decreases monotonically in the range $21<\alpha<$ 2, with $\nu(\alpha=2)\simeq 1.4$, and then grows again to $\nu\simeq 2.0$ for $\alpha=2.7$ (see Supplemental Material).
We conclude by discussing the behavior of the single-particle density matrix $\mathcal{G}(\ell)=\langle\,{b}^{\dagger}_{i}{b}_{i+\ell}\rangle$ in the superfluid phase as a function of distance $\ell$ near the critical point. Example results are shown in Fig. 5(a-d) as a function of the chord distance $c(\ell)=\sin(\pi\ell/L)$ to account for periodic boundary conditions. In the figure, we present results for Eq. (1) for hard-core bosons ($t/U\rightarrow 0$) at half-filling, which allows us to explore the behavior of $\mathcal{G}(\ell)$ up to sizes as large as $L=1024$ and to directly compare to existing predictions for the long-range XY model [35, 44]. Figure 5 shows that $\mathcal{G}(\ell)$ decays with increasing $c(\ell)$ up to the largest sizes considered, for all $1<\alpha\leq 3$. Interestingly, for each $\alpha$ and given size $L$, this large-distance decay is very well approximated by a simple power-law as $\mathcal{G}(l)\sim c(l)^{-\gamma}$ (see Supplemental Material). However, in contrast to familiar 1D quantum liquids with short-range hopping, the value of $\gamma$ is $L$-dependent and decreases with $L$ (see also Insets) raising the question of its asymptotic value in the thermodynamic limit. The fitted values of $\gamma$ are plotted vs $1/L$ in panel (e), showing values consistent with $\gamma=0$ for $1<\alpha\lesssim 2$ and finite $\gamma>0$ for $2\lesssim\alpha\leq 3$ in the thermodynamic limit. For $1<\alpha\lesssim 2$, this is a clear sign of long-range order, a result consistent with literature for related models [18, 35, 45]. For $2\lesssim\alpha\leq 3$, our results can be also fitted by a combination of the power-law decay and the $L$-dependent constant (see Supplemental Material). The former fit is better by eliminating one fitting parameter and produces cleaner results; given very small values of $\gamma$ observed we term the region $2\lesssim\alpha\leq 3$ as anomalous quasi-long-range-order, departing from expectations in literature. The pure power-law analysis is further corroborated by revealing finite-size scaling of the density matrix, $\mathcal{G}(\ell)$, at the largest scale $\ell=L/2$ in Fig. 5(f). It represents the condensate fraction in the system with finite condensate, and, thus, is expected to saturate to an $L$-independent value in the presence of long-range order. Otherwise, it decays as a power-law with increasing $L$ in the presence of quasi-long-range order. Figure 5(f) shows that $\mathcal{G}(L/2)$ saturates to a constant for $\alpha<2$ and decays as a power-law with increasing $L$ for $\alpha>2$. We find no evidence to support a speculation that this behavior will change at length scales larger than what were computed here. In the Supplemental Material, we provide additional data for $\mathcal{G}(\ell)$ in the superfluid phase of soft-core bosons, computed for distances up to $L=512$. These results demonstrate correlation behavior consistent with that shown in Fig. 5. Additionally, we present data for $\mathcal{G}(\ell)$ in the Mott insulating (MI) phase, where it exhibits power-law decay. In this phase, the decay exponent $\gamma$ converges to the power-law hopping exponent $\alpha$ in the thermodynamic limit.
[…]
Our predictions for the superfluid dispersion relation and the correlation functions can be directly measured in experiments with dipolar atoms and molecules ($\alpha=3$, [15, 47, 48, 49, 50]) and cold ions ($1<\alpha\lesssim 3$, [16, 22, 23]). Our work provides exact results to benchmark theories for long-range quantum models and opens up multiple other research directions, including the nature of the groundstate in higher dimensions.
Acknowledgments: T.G. and G.P. gratefully acknowledge discussions with Guido Giachetti. This research has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie project 955479 (MOQS), the Horizon Europe programme HORIZON-CL4-2021-DIGITAL-EMERGING-01-30 via the project 101070144 (EuRyQa) and from the French National Research Agency under the Investments of the Future Program projects ANR-21-ESRE-0032 (aQCess), ANR-22-CE47-0013-02 (CLIMAQS) and ANR-23-CE30- 50022-02 (SIX). NP acknowledges support from the National Science Foundation under Grant No. DMR2335904.

### v2

@@
 In the figures, the values of $\langle\mathcal{W}^{2}\rangle_{L_{\text{max}}}$ for the largest sizes $L_{\text{max}}$ used in the computations have been subtracted for clarity.
-Panels (a) and (b) show a clear crossing of $\langle\mathcal{W}^{2}\rangle$ vs $t/U$ when plotted for different $L$ at values $(t/U)_{c}=0.043\pm 0.0005$ and $(t/U)_{c}=0.131\pm 0.0008$, respectively.
+Panels (a) and (b) show a clear crossing of $\langle\mathcal{W}^{2}\rangle$ vs $t/U$ when plotted for different $L$ at values $(t/U)_{c}=0.0430\pm 0.0005$ and $(t/U)_{c}=0.131\pm 0.001$, respectively.
 These crossings correspond to a quantum phase transition at the respective values of the critical ratio $(t/U)_{c}$ (see Insets for further extrapolation to thermodynamic limit).
@@
 […]
-Our predictions for the superfluid dispersion relation and the correlation functions can be directly measured in experiments with dipolar atoms and molecules ($\alpha=3$, [15, 47, 48, 49, 50]) and cold ions ($1<\alpha\lesssim 3$, [16, 22, 23]).
+Our predictions for the superfluid dispersion relation and the correlation functions can be directly measured in experiments with dipolar atoms and molecules ($\alpha=3$, [15, 47, 48, 49, 50]) as well as cold ions ($1<\alpha\lesssim 3$, [16, 22, 23]) and are directly relevant to recent experiments with subradiant dipolar excitons in double quantum well systems [51].
 Our work provides exact results to benchmark theories for long-range quantum models and opens up multiple other research directions, including the nature of the groundstate in higher dimensions.